hhanan:

The population in Smalltown in 2010 was 45,230 people and is growing exponentially at a rate of 2%. What will the population be in 2030?

XioGonz:

P = population T= years that followed 2010 = 45230e0.02t Population in 2030 = P(20) =45230e^0.02T-(20) = 45230e^0.4 = 67475 @jhonyy9

hhanan:

thank you

jhonyy9:

good job @XioGonz